# Entropy Functions on Two-Dimensional Faces of Polymatroidal Region of Degree Four: Part II: Information Theoretic Constraints Breed New Combinatorial Structures

### v1

@@ Abstract
Characterization of entropy functions is of fundamental importance in information theory. By imposing constraints on their Shannon outer bound, i.e., the polymatroidal region, one obtains the faces of the region and entropy functions on them with special structures. In this series of two papers, we characterize entropy functions on the $2$-dimensional faces of the polymatroidal region $\Gamma_{4}$. In Part I, we formulated the problem, enumerated all $59$ types of $2$-dimensional faces of $\Gamma_{4}$ by an algorithm, and fully characterized entropy functions on $49$ types of them. In this paper, i.e., Part II, we will characterize entropy functions on the remaining $10$ types of faces, among which $8$ types are fully characterized and $2$ types are partially characterized. To characterize these types of faces, we introduce some new combinatorial design structures which are interesting themself.
@@ I Introduction
The region in $\mathcal{H}_{n}$ bounded by such inequalities, denoted by $\Gamma_{n}$, is called the polymatroidal region of degree $n$. Thus, $\Gamma_{n}$ is an outer bound on $\Gamma^{*}_{n}$. For more about entropy functions, we referred the readers to [1, Chapter 13-15].
Traditionally, entropy functions are characterized by information inequalities. Those inequalities derived by polymatroidal axioms are called Shannon-type, as they correspond to the non-negativity of Shannon information measures. Since 1998, a series of non-Shannon-type information inequalities, among which Zhang-Yeung inequality is the first one [2], were discovered [3][4][5]. Thus $\overline{\Gamma^{*}_{n}}$, the closure of $\Gamma^{*}_{n}$, is strictly included in $\Gamma_{n}$ when $n\geq 4$. Each information inequality determines an outer bound on $\Gamma^{*}_{n}$, as those set functions in $\mathcal{H}_{n}$ dissatisfy it must be located outside $\Gamma^{*}_{n}$. In this series of two papers, we develop a system of entropy function characterization from the perspective of faces of $\Gamma_{n}$, which covers the traditional inequality characterization.
@@
Many information-theoretic problems can be considered as applications of entropy function characterizations on faces of $\Gamma_{n}$. In a series of three papers [13, 14, 15], Matúš and Studneý solved the probabilistic conditional independence problem for four random variables. Note each class of conditional independence constraints, which is called a semimatroid in their papers, determines a face $F$ of $\Gamma_{n}$. The solution to this problem, i.e., the probabilistic-representability of a semimatroid determines whether the relative interior of the corresponding face $F$ interset with $\Gamma^{*}_{n}$. This problem thus can be considered as partial characterizations of the faces of $\Gamma_{n}$. In [16], Yan, Yeung, and Zhang proved a formula involving $\Gamma_{n}^{*}$ for the capacity of multi-source multi-sink network coding. Those constraints in the formula induced by network topology and source independency form a face of $\Gamma_{n}$, which shows that this holy-grail network coding problem corresponds to the entropy function characterizations on this face. For the secret-sharing problem, see [17] for example, the perfect secrecy and decoding correctness conditions of an access structure determine a face $F$ on $\Gamma_{n}$, and the information ratio of the secret-sharing problem can be considered as an optimization problem whose feasible region is $F\cap\Gamma^{*}_{n}$. Other problems, such as distributed data storage [18], coded caching [19], Markov random fields [20], and relational database [21] are also related to the entropy regions on the faces of $\Gamma_{n}$.
@@
Though the information theory problems discussed above usually involve more than four random variables, and the corresponding faces are of dimension higher than $2$, following the characterizations of extreme rays. i.e., $1$-dimensional faces containing matroid [6][7][8], and $2$-dimensional faces of $\Gamma_{3}$ [9][10], in this series of two papers, we characterize entropy functions on the $2$-dimensional faces of $\Gamma_{4}$, which may serve as stepping stones to the general cases of this problem. In Part I [22], we enumerated all $59$ types $2$-dimensional faces of $\Gamma_{4}$ by an algorithm and completely characterized $49$ types of them. In this part, we characterize the remaining $10$ types of faces, among which $8$ types are fully characterized and $2$ types are partially characterized. To characterize these types of faces, we adopt two sets of combinatorial structures, that is, mixed-level variable-strength orthogonal arrays and orthogonal Latin hypercubes, and the characterizations breed some new combinatorial design structures which are interesting themself.
@@
The rest of this paper is organized as follows. In Section II, for self-contain, we give the preliminaries on integer polymatroids and matroids, polyhedral cones, mixed-level variable-strength orthogonal arrays and its relationship to orthogonal Latin hypercubes. We then list the results of characterization of extreme rays of $\Gamma_{4}$ which has been done in Part I without proofs. In Section III, we complete the characterization of the remaining $10$ types of $2$-dimensional faces of $\Gamma_{4}$, and the results are summarized in Table LABEL:table:2. The correspondence between the faces and the theorems to characterizing them in two papers are summerized in Table 4.
@@ II-B Faces of a polyhedral cone
Let $C\subseteq\mathbb{R}^{d}$ be a full-dimensional polyhedral cone. For a hyperplane $P$ containing the origin $O$ in $\mathbb{R}^{d}$, if $C\subseteq P^{+}$, where $P^{+}$ is one of the two halfspaces corresponding to $P$, $F\triangleq C\cap P$ is called a (proper) face of $C$, while $C$ itself is its improper face. When $\dim F=d-1$, $F$ is called a facet of $C$, and when $\dim F=1$, $F$ is an extreme ray of $C$. Either the set of all facets or the set of all extreme rays of $C$ uniquely determines the cone, and they are called H-representation and $V$-representation of the cone, respectively. For each face $F$ of the cone, it can be written as the intersection of the facets of the cone that contains $F$, or the convex hull of the extreme rays contained in $F$. We also call them the H-representation and V-representation of $F$, respectively. More details on this topic are referred [24].
As we discussed in Section I, $\Gamma_{n}$ is a polyhedral cone in $\mathcal{H}_{n}$ determined by polymatroidal axioms in (I.1)-(I.3), They are equivalent to the following elemental inequalities
@@
for $\mathcal{W}^{\alpha}_{2}$ with $\alpha\subseteq N_{4}$, $|\alpha|=2$, it is called a wheel matroid with order 22, [^1] We adopt the notation and terminology in [23, Section 8.4] and it is a matroid with two parallel elements in $\alpha$, and each element in $\alpha$ and the other two elements in $N_{4}$ form a $U_{2,3}$;
@@
for $\hat{U}^{i}_{2,5}$ with $i\in N_{4}$, it is a polymatroid whose free expansion is a $U_{2,5}$, [^2] See [26, Theorem 1.3.6] and [27, Theorem 4] for the definition of free expansion. and its rank function is defined by
@@ II-C Mixed-level variable-strength orthogonal arrays and orthogonal Latin hypercubes
To characterize entropy functions on the extreme rays of $\Gamma_{4}$, we adopt two sets of combinatorial structures, that is, mixed-level variable-strength orthogonal arrays and orthogonal Latin hypercubes, which are equivalent in some sense but each have their own advantages. For mixed-level variable-strength orthogonal arrays, it will be easier to be generalized to the cases with more random variables and higher dimensions, and the symmetries between random variables are more straightforward. In contrast, the language of orthogonal Latin hypercubes will be better visualized for the cases for four random variables. Furthermore, they will also play important roles in characterizing $2$-dimensional faces of $\Gamma_{4}$, especially those with two extreme rays containing integer polymatroids of rank exceeding $1$.
@@
Given an integer polymatroid $P=(N,\mathbf{r})$ with $\mathbf{r}(N)\geq 2$ and an integer $v>1$, a $v^{\mathbf{r}(N)}\times n$ array $\mathbf{T}$ with columns indexed by $N$, whose entries of a column indexed by $i\in N$ are from a set $I_{i}$ with cardinality $v^{\mathbf{r}(i)}-1$, is called a mixed-level variable-strength orthogonal array(MVOA) induced by $P$ with base level $v$ if for each $A\subseteq N$, $v^{\mathbf{r}(N)}\times|A|$ subarray $\mathbf{T}(A)$ of $\mathbf{T}$ consisting of columns indexed by $A$ satisfies the following condition: each row of $\mathbf{T}(A)$ occurs in $\mathbf{T}(A)$ exactly $v^{\mathbf{r}(N)-\mathbf{r}(A)}$ times. We also call such $\mathbf{T}$ a $\mathrm{MVOA}(P,v)$.
@@
Usually, we set $I_{i}=\mathbb{I}_{v^{\mathbf{r}(i)}-1}$, unless otherwise specified, where $\mathbb{I}_{u}\triangleq\{0,1,\ldots,u-1\}$ for any integer $u\geq 1$. When the integer polymatroid is a matroid $M$, its induced MVOA is reduced to a variable-strength orthogonal array $\mathrm{VOA}(M,v)$. When a matroid is a uniform matroid $U_{t,n}$, the induced $\mathrm{VOA}(U_{t,n},v)$ is an orthogonal array with index unity, usually denoted by $\mathrm{OA}(t,n,v)$ [6, 7].
@@
To discribe MVOAs, we introduce some notations which will be used throughout this paper. The $v^{\mathbf{r}(N_{n})}\times|A|$ subarray of $\mathbf{T}$ consisting of columns indexed by $A$ is denoted by $\mathbf{T}(A)$. For simplicity, we drop the brackets of $\mathbf{T}(A)$, e.g., we write $\mathbf{T}(\{1,2,3\})$ as $\mathbf{T}(1,2,3)$. Let $\mathbf{T}(A;j)$ denote the $j$-th row of $\mathbf{T}(A)$.
@@
For the orthogonal Latin hypercubes, we will not give a general defintion in this paper. As on the one hand, for some extreme rays containing an integer polymatroid other than uniform matroid, the variants of traditional orthogonal Latin hypercubes are utilized. On the other hand, in this paper, we only consider the cases corresponding to integer polymatroids in $\Gamma_{4}$, case by case discussion is enough. For the general form of these variants, it can be considered in the future research.
@@
We begin our discussion from Latin square, the non trivial orthogonal Latin hypercubes with the smallest parameters, which corresponding to the smallest connected matroid with rank exceeding $1$. For uniform matroid $U_{2,3}$, $\mathrm{VOA}(U_{2,3},v)$ corresponds to a Latin square $S$ of order $v$ with row and column indices and symbols in $\mathbb{I}_{v}$, that is, a $v\times v$ square $S$ with indexed by $i,j\in\mathbb{I}_{v}$ with symbols in the cells of each row $\{S(i,j),j\in\mathbb{I}_{v}\}$ all distinct. For each row $(i,j,k)$ of a $\mathrm{VOA}(U_{2,3},v)$ $\mathbf{T}$, it corresponds to the cell $S(i,j)$ with symbol $k$. Here we also define Latin square of the zeroth-kind and the second-kind  [^3] The definition of Latin squares of the zeroth-kind and the second-kind are in the spirit of the definition of the “Latin cubes of first order and second order” in [28]. As usually “order” used for the size of a Latin square, in our paper, we use “kind” instead. In the following, Latin cubes of the first kind and second kind will be defined similarly. which are trivial themselve, but can be considered as building blocks in construction of other combinatorial structure in this paper, while for a traditional Latin square, we call it of the first kind. For a $v\times v$ square $S$ in all cells a unique symbol 0, we call it a Latin square of the zeroth kind. It corresponds to a $\mathrm{VOA}(U^{1,2;3}_{1,1},v)$, where $U_{1,1}^{1,2;3}$ is a matroid whose rank function is the summaton of the rank functions of $U^{1;3}_{1,1}$ and $U^{2;3}_{1,1}$. For a $v\times v$ square $S$, in cells all symbols distinct from $\mathbb{I}_{v^{2}}$, we call it a Latin square of the second kind. It corresponds to a $\mathrm{MVOA}(U^{13,23;3}_{1,2},v)$, where $U^{13,23;3}_{1,2}$ is an integer polymatroid whose rank function is the summation of the rank functions of $U^{13;3}_{1,2}$ and $U^{23;3}_{1,2}$. For a traditional Latin square, or a Latin square of the first kind, we call it a Latin square and omit “of the first kind” if there is no ambiguity.
@@
The following are Latin squares of the zeroth kind, the first kind, and the second kind, and their corresponding MVOAs.
@@
  • For a $\mathrm{VOA}(U^{123}_{2,3},v)$ $\mathbf{T}$, $\mathbf{T}(1,2,3)$ forms a $\mathrm{VOA}(U_{2,3},v)$, and all the entries of the column $\mathbf{T}(4)$ can be a constant symbol 0. It corresponds to the a pair of Latin squares sharing the same row and column indices and one of the first-kind and the other of the zeroth-kind, where $\mathbf{T}(1)$ and $\mathbf{T}(2)$, $\mathbf{T}(3)$, $\mathbf{T}(4)$ corresponds to the indices of rows and columns, the symbols of the first square and symbols of the second square, respectively.
@@
[…]
  • For a $\mathrm{MVOA}(\hat{U}^{4}_{2,5},v)$ $\mathbf{T}$, it corresponds to a pair of $v\times v$ squares $(S_{1},S_{2})$, where, $S_{1}$ is a Latin square of the first kind, and $S_{2}$ is a Latin square of the second kind.
@@
The following are the orthogonal Latin hypercubes and their corresponding MOVAs discussed above.

### v2

@@ Abstract
The characterization of entropy functions is of fundamental importance in information theory. By imposing constraints on their Shannon outer bound, i.e., the polymatroidal region, one obtains the faces of the region and entropy functions on them with special structures. In this series of two papers, we characterize entropy functions on the $2$-dimensional faces of the polymatroidal region $\Gamma_{4}$. In Part I, we formulated the problem, enumerated all $59$ types of $2$-dimensional faces of $\Gamma_{4}$ by an algorithm, and fully characterized entropy functions on $49$ types of them. In this paper, i.e., Part II, we will characterize entropy functions on the remaining $10$ types of faces, among which $8$ types are fully characterized, and $2$ types are partially characterized. To characterize these types of faces, we introduce some new combinatorial design structures that are interesting in themselves.
@@ I Introduction
The region in $\mathcal{H}_{n}$ bounded by such inequalities, denoted by $\Gamma_{n}$ is called the polymatroidal region of degree $n$. Thus, $\Gamma_{n}$ is an outer bound on $\Gamma^{*}_{n}$. For more about entropy functions, we referred the readers to [1, Chapter 13-15].
Traditionally, entropy functions are characterized by information inequalities. Those inequalities derived by polymatroidal axioms are called Shannon-type, as they correspond to the non-negativity of Shannon information measures. Since 1998, a series of non-Shannon-type information inequalities, among which Zhang-Yeung inequality is the first one [2], were discovered [3][4][5]. Thus $\overline{\Gamma^{*}_{n}}$, the closure of $\Gamma^{*}_{n}$, is strictly included in $\Gamma_{n}$ when $n\geq 4$. In 2007, Matúš showed that for $n\geq 4$, there exist infinitely non-redundant non-Shannon-type inequalities, indicating that $\overline{\Gamma_{n}^{*}}$ is not polyhedral [6]. Each information inequality determines an outer bound on $\Gamma^{*}_{n}$, as those set functions in $\mathcal{H}_{n}$ that dissatisfy it must be located outside $\Gamma^{*}_{n}$. In this series of two papers, we develop a system of entropy function characterization from the perspective of faces of $\Gamma_{n}$, which covers the traditional inequality characterization.
@@
Many information-theoretic problems can be considered as applications of entropy function characterizations on faces of $\Gamma_{n}$. In a series of three papers [14, 15, 16], Matúš and Studneý solved the probabilistic conditional independence problem for four random variables. Note each class of conditional independence constraints, which is called a semimatroid in their papers, determines a face $F$ of $\Gamma_{n}$. The solution to this problem, i.e., the probabilistic-representability of a semimatroid determines whether the relative interior of the corresponding face $F$ intersects with $\Gamma^{*}_{n}$. This problem thus can be considered as partial characterizations of the faces of $\Gamma_{n}$. In [17], Yan, Yeung, and Zhang proved a formula involving $\Gamma_{n}^{*}$ for the capacity of multi-source multi-sink network coding. Those constraints in the formula induced by network topology and source independency form a face of $\Gamma_{n}$, which shows that this holy-grail network coding problem corresponds to the entropy function characterizations on this face. For the secret-sharing problem, see [18] for example, the perfect secrecy and decoding correctness conditions of an access structure determine a face $F$ on $\Gamma_{n}$, and the information ratio of the secret-sharing problem can be considered as an optimization problem whose feasible region is $F\cap\Gamma^{*}_{n}$. Other problems, such as distributed data storage [19], coded caching [20], Markov random fields [21], and relational database [22] are also related to the entropy regions on the faces of $\Gamma_{n}$.
@@
Though the information theory problems discussed above usually involve more than four random variables, and the corresponding faces are of dimension higher than $2$, following the characterizations of extreme rays. i.e., $1$-dimensional faces containing matroid [7][8][9], and $2$-dimensional faces of $\Gamma_{3}$ [10][11], in this series of two papers, we characterize entropy functions on the $2$-dimensional faces of $\Gamma_{4}$, which may serve as stepping stones to the general cases of this problem. In Part I [23], we enumerated all $59$ types $2$-dimensional faces of $\Gamma_{4}$ by an algorithm and completely characterized $49$ types of them. In this part, we characterize the remaining $10$ types of faces, among which $8$ types are fully characterized and $2$ types are partially characterized. To characterize these types of faces, we adopt two sets of combinatorial structures, that is, mixed-level variable-strength orthogonal arrays and orthogonal Latin hypercubes, and the characterizations breed some new combinatorial design structures that are interesting in themselves.
@@
The rest of this paper is organized as follows. In Section II, for self-contained, we give the preliminaries on integer polymatroids and matroids, polyhedral cones, mixed-level variable-strength orthogonal arrays and their relationship to orthogonal Latin hypercubes. We then list the results of the characterization of extreme rays of $\Gamma_{4}$, which has been done in Part I, without proofs. In Section III, we complete the characterization of the remaining $10$ types of $2$-dimensional faces of $\Gamma_{4}$, and the results are summarized in Table LABEL:table:2 together with those in Part I. The correspondences between the faces and the theorems characterizing them in two papers are summarized in Table 4.
@@ II-B Faces of a polyhedral cone
Let $C\subseteq\mathbb{R}^{d}$ be a full-dimensional polyhedral cone. For a hyperplane $P$ containing the origin $O$ in $\mathbb{R}^{d}$, if $C\subseteq P^{+}$, where $P^{+}$ is one of the two halfspaces corresponding to $P$, $F\triangleq C\cap P$ is called a (proper) face of $C$, while $C$ itself is its improper face. When $\dim F=d-1$, $F$ is called a facet of $C$, and when $\dim F=1$, $F$ is an extreme ray of $C$. Either the set of all facets or the set of all extreme rays of $C$ uniquely determines the cone, and they are called the H-representation and $V$-representation of the cone, respectively. For each face $F$ of the cone, it can be written as the intersection of the facets of the cone that contain $F$, or the convex hull of the extreme rays contained in $F$. We also call them the H-representation and V-representation of $F$, respectively. More details on this topic are referred to [25].
As we discussed in Section I, $\Gamma_{n}$ is a polyhedral cone in $\mathcal{H}_{n}$ determined by polymatroidal axioms in (I.1)-(I.3). They are equivalent to the following elemental inequalities
@@
for $\mathcal{W}^{\alpha}_{2}$ with $\alpha\subseteq N_{4}$, $|\alpha|=2$, it is called a wheel matroid with order 22, [^1] We adopt the notation and terminology in [24, Section 8.4] and it is a matroid with two parallel elements in $\alpha$, and each element in $\alpha$ and the other two elements in $N_{4}$ form a $U_{2,3}$;
@@
for $\hat{U}^{i}_{2,5}$ with $i\in N_{4}$, it is a polymatroid whose free expansion is a $U_{2,5}$, [^2] See [27, Theorem 1.3.6] and [28, Theorem 4] for the definition of free expansion. and its rank function is defined by
@@ II-C Mixed-level variable-strength orthogonal arrays and orthogonal Latin hypercubes
To characterize entropy functions on the extreme rays of $\Gamma_{4}$, we adopt two sets of combinatorial structures, that is, mixed-level variable-strength orthogonal arrays and orthogonal Latin hypercubes, which are equivalent in some sense, but each has its own advantages. For mixed-level variable-strength orthogonal arrays, it will be generalized more easily to the cases with more random variables and higher dimensions, and the symmetries between random variables are more straightforward. In contrast, the language of orthogonal Latin hypercubes will be better visualized for the cases of four random variables. Furthermore, they will also play important roles in characterizing $2$-dimensional faces of $\Gamma_{4}$, especially those with two extreme rays containing integer polymatroids of rank exceeding $1$.
@@
Given an integer polymatroid $P=(N_{n},\mathbf{r})$ with $\mathbf{r}(N_{n})\geq 2$ and an integer $v>1$, a $v^{\mathbf{r}(N_{n})}\times n$ array $\mathbf{T}$ with columns indexed by $N_{n}$, whose entries of a column indexed by $i\in N_{n}$ are from a set $I_{i}$ with cardinality $v^{\mathbf{r}(i)}-1$, is called a mixed-level variable-strength orthogonal array(MVOA) induced by $P$ with base level $v$ if for each $A\subseteq N_{n}$, $v^{\mathbf{r}(N_{n})}\times|A|$ subarray $\mathbf{T}(A)$ of $\mathbf{T}$ consisting of columns indexed by $A$ satisfies the following condition: each row of $\mathbf{T}(A)$ occurs in $\mathbf{T}(A)$ exactly $v^{\mathbf{r}(N_{n})-\mathbf{r}(A)}$ times. We also call such $\mathbf{T}$ a $\mathrm{MVOA}(P,v)$.
@@
Usually, we set $I_{i}=\mathbb{I}_{v^{\mathbf{r}(i)}-1}$, unless otherwise specified, where $\mathbb{I}_{u}\triangleq\{0,1,\ldots,u-1\}$ for any integer $u\geq 1$. When the integer polymatroid is a matroid $M$, its induced MVOA is reduced to a variable-strength orthogonal array $\mathrm{VOA}(M,v)$[7],[8],[9]. When a matroid is a uniform matroid $U_{t,n}$, the induced $\mathrm{VOA}(U_{t,n},v)$ is an orthogonal array with index unity, usually denoted by $\mathrm{OA}(t,n,v)$ [7, 8].
@@
To describe MVOAs, we introduce some notations which will be used throughout this paper. The $v^{\mathbf{r}(N_{n})}\times|A|$ subarray of $\mathbf{T}$ consisting of columns indexed by $A$ is denoted by $\mathbf{T}(A)$. For simplicity, we drop the brackets of $\mathbf{T}(A)$, e.g., we write $\mathbf{T}(\{1,2,3\})$ as $\mathbf{T}(1,2,3)$. Let $\mathbf{T}(A;j)$ denote the $j$-th row of $\mathbf{T}(A)$.
@@
For the orthogonal Latin hypercubes, we will not give a general definition in this paper. Because, on the one hand, for some extreme rays containing an integer polymatroid other than a uniform matroid, the variants of traditional orthogonal Latin hypercubes are utilized. On the other hand, in this paper, we only consider the cases corresponding to integer polymatroids in $\Gamma_{4}$, and so, a case-by-case discussion is enough. For the general form of these variants, it can be studied in future research.
@@
We begin our discussion from Latin squares, the non-trivial orthogonal Latin hypercubes with the smallest parameters, which corresponds to the smallest connected matroid with rank exceeding $1$. For uniform matroid $U_{2,3}$, a $\mathrm{VOA}(U_{2,3},v)$ corresponds to a Latin square $S$ of order $v$ with row and column indices and symbols in $\mathbb{I}_{v}$, that is, a $v\times v$ square $S$ with cells indexed by $(i,j)\in\mathbb{I}^{2}_{v}$ and symbols in the cells of each row $\{S(i,j),j\in\mathbb{I}_{v}\}$ and each column $\{S(i,j),i\in\mathbb{I}_{v}\}$ all distinct. For each row $(i,j,k)$ of a $\mathrm{VOA}(U_{2,3},v)$ $\mathbf{T}$, it corresponds to the cell $S(i,j)$ with symbol $k$. Here, we also define Latin squares of the zeroth-kind and the second-kind, [^3] The definition of Latin squares of the zeroth-kind and the second-kind are in the spirit of the definition of the “Latin cubes of first order and second order” in [29]. “Order” is usually used for the size of a Latin square, so in our paper, we use “kind” instead. In the following, Latin cubes of the first kind and the second kind will be defined similarly. which are trivial themselves, but can be considered as building blocks in construction of other combinatorial structures in this paper, while for a traditional Latin square, we call it of the first kind. For a $v\times v$ square $S$ with symbols in all cells, each identical to 0, we call it a Latin square of the zeroth kind. It corresponds to a $\mathrm{VOA}(U^{1,2;3}_{1,1},v)$, where $U_{1,1}^{1,2;3}$ is a matroid whose rank function is the summation of the rank functions of $U^{1;3}_{1,1}$ and $U^{2;3}_{1,1}$. For a $v\times v$ square $S$ with symbols in all cells distinct from $\mathbb{I}_{v^{2}}$, we call it a Latin square of the second kind. It corresponds to a $\mathrm{MVOA}(U^{13,23;3}_{1,2},v)$, where $U^{13,23;3}_{1,2}$ is an integer polymatroid whose rank function is the summation of the rank functions of $U^{13;3}_{1,2}$ and $U^{23;3}_{1,2}$. For a traditional Latin square, or a Latin square of the first kind, we call it a Latin square and omit “of the first kind” if there is no ambiguity.
@@
In Table 2, we list Latin squares of the zeroth kind, the first kind, and the second kind, and their corresponding MVOAs for $v=3$.
@@
  • For a $\mathrm{VOA}(U^{123}_{2,3},v)$ $\mathbf{T}$, $\mathbf{T}(1,2,3)$ forms a $\mathrm{VOA}(U_{2,3},v)$, and all the entries of the column $\mathbf{T}(4)$ can be a constant symbol 0. It corresponds to a pair of Latin squares sharing the same row and column indices and one of which is the first-kind, and the other of which is the zeroth-kind, where $\mathbf{T}(1)$ and $\mathbf{T}(2)$, $\mathbf{T}(3)$, $\mathbf{T}(4)$ corresponds to the indices of rows and columns, the symbols of the first square and symbols of the second square, respectively.
@@
  • For a $\mathrm{VOA}(U_{3,4},v)$ $\mathbf{T}$, it corresponds to a Latin cube of order $v$ (type 2 and the first kind, permutation cube [30, VI.22.33]), where entries in $\mathbf{T}(1)$, $\mathbf{T}(2)$ and $\mathbf{T}(3)$ correspond to the indices of three dimensions, respectively, and $\mathbf{T}(4)$ corresponds to the symbols. A Latin cube of order $v$ and the first kind is a $v\times v\times v$ cube with symbols in each line $\{C(i,j,k),i\in\mathbb{I}_{v}\}$ for fixed $j,k\in\mathbb{I}_{v}$, $\{C(i,j,k),j\in\mathbb{I}_{v}\}$ for fixed $i,k\in\mathbb{I}_{v}$, and $\{C(i,j,k),i\in\mathbb{I}_{v}\}$ for fixed $i,j\in\mathbb{I}_{v}$ all distinct.
  • For a $\mathrm{MVOA}(\hat{U}^{4}_{2,5},v)$ $\mathbf{T}$, it corresponds to a pair of $v\times v$ squares $(S_{1},S_{2})$, where $S_{1}$ is a Latin square of the first kind, and $S_{2}$ is a Latin square of the second kind.
@@
In Table 3, we list orthogonal Latin hypercubes and their corresponding MVOAs discussed above for $v=3$.